# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 1 Introduction
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
determines a weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
$P_{d}(\overline{v},\,n)=h(W(v,\,n))\,.$
See Section 5 for the importance of (the general form of) this equality.
We structure our article as follows. In Section 2 we review infinite series with complex summands and the domain $\mathbb{C}[[x]]$ of generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [17]; see Section 5 for more details on [17].
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [3, 8, 10, 12, 22, 23, 28] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop this method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series and their coefficients are determined by finite operations. We also use formal power series, but our extension allows operations based on limit transitions. See [15] for more applications of the semiformal method.
@@ 2 Series and generating functions in $\mathbb{C}$
We review series and generating functions in $\mathbb{C}$. A series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@
    If the set $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$ the proper limit
@@
(In fact, the independence of $s$ on $f$ may be omitted from the formulation.) We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$. If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
@@
is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $F_{U}(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
@@
where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. Thus for any $U\in H_{1}$ the sums form the function $F_{U}\colon\mathbb{C}_{<1}\to\mathbb{C}$.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. A walk $w$ in $K_{\mathbb{N}}$ is any $(n+1)$-tuple $w$ of vertices $v_{i}\in\mathbb{N}$,
$w=\langle v_{0},\,v_{1},\,\dots,\,v_{n}\rangle\,$,
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length of $w$ is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
$h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ with length $n$ such that $v_{0}=1$.
Definition 2.1 (lightness)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We call the weight $h$ light if for every $n\in\mathbb{N}_{0}$, the series
$\sum_{w\in W(n)}h(w)$}
@@
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges. However, we cannot compare the sizes of sums
$\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)$}
— we call it the subseries bound principle.
If $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and $\alpha,\beta\in\mathbb{C}$, the series

### v8

@@ 1.2 Pólya’s theorem for $\overline{v}=\overline{0}$ via the symbolic method
Now we derive (1) for $\overline{v}=\overline{0}$ by the symbolic method. The grid graph $G_{d}$ is the $2d$-regular graph $G_{d}=\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|\overline{a}_{i}-\overline{b}_{i}|=1\,.$}
We take the constant map $h\colon E_{d}\to\{\frac{1}{2d}\}$ and extend it to walks in $G_{d}$ by $h(w)=(2d)^{-n}$ where $n$ ($\in\mathbb{N}_{0}$) is the length of the walk $w$. Let $A$ be the set of (finite) walks in $G_{d}$ that start at the origin $\overline{0}$ and later revisit it. The size function $s$ is defined on every walk $w$ in $G_{d}$ and $s(w)$ is the length of $w$. Then for every $d\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$,
$P_{d}(\overline{0},\,n)=\sum_{w\in A_{n}}h(w)=:h(A_{n})$}
(recall that $A_{n}=\{w\in A\colon\;s(w)=n\}$). We are interested in the main generating function
$A_{h}(x)=\sum_{n\geq 0}h(A_{n})x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
We obtain a formula for $\lim_{n\to\infty}h(A_{n})$ with the help of three auxiliary generating functions:
$B_{h}(x)=\sum_{n\geq 0}h(B_{n})x^{n},\ C_{h}(x)=\sum_{n\geq 0}h(C_{n})x^{n},\ D_{h}(x)=\sum_{n\geq 0}h(D_{n})x^{n}\,.$}
Here $B$ is the set of closed walks in $G_{d}$ that begin and end at $\overline{0}$, $C\subset B$ is the subset of closed walks with inner vertices different from $\overline{0}$ and $D$ is the set of all walks in $G_{d}$ starting at $\overline{0}$. We have $h(A_{0})=h(C_{0})=0$ and $h(B_{0})=h(D_{0})=1$.
The first formal phase of the symbolic method yields two formal identities,
$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}\,\text{ and }\,C_{h}(x)=1-\frac{1}{B_{h}(x)}\ \ (\in\mathbb{C}[[x]])\,.$}
The second equality reflects the fact that $h(D_{n})=1$ for every $n\in\mathbb{N}_{0}$. The other two equalities are also easy to prove; in Proposition 3.1 we prove them in a much more general setting. The first identity implies that for every $n\in\mathbb{N}_{0}$,
$h(A_{n})=\sum_{j=0}^{n}h(C_{j})\,.$}
We proceed to the second analytic phase. We consider the sums
$C_{h}(1)=\sum_{n\geq 0}h(C_{n}):=\lim_{n\to\infty}\sum_{j=0}^{n}h(C_{j})\ \ (\in\mathbb{C})$}
and $B_{h}(1)$ which is defined similarly. Since the formal power series $B_{h}(x)$ and $C_{h}(x)$ have coefficients in $[0,1]$, the formal identity $C_{h}(x)=1-\frac{1}{B_{h}(x)}$ in $\mathbb{C}[[x]]$ converts to the identity
$C_{h}(1)=1-\frac{1}{B_{h}(1)}\ \ (\in[0,\,1])\,$,}
where $B_{h}(1)\in[1,\,+\infty)\cup\{+\infty\}$. We rigorously prove it. We formally have $B_{h}(x)C_{h}(x)=B_{h}(x)-1$. Since $0\leq h(B_{n}),h(C_{n})\leq h(D_{n})=1$ for every $n\in\mathbb{N}_{0}$, the formal power series $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for every $x\in[0,1)$. Using the operations of linear combination, grouping, and product for complex series in Propositions 2.2, 2.4, and 2.5, respectively, we get that
$\text{$B_{h}(x)C_{h}(x)=B_{h}(x)-1$ for every number $x\in[0,\,1)$}\,.$
Since $B_{h}(x)\geq 1$ for every $x\in[0,1)$, we can divide by it and get that
$\text{$C_{h}(x)=1-\frac{1}{B_{h}(x)}$ for every $x\in[0,\,1)$}\,.$}
Using the arithmetic of functional limits and Theorem 2.18 (Abel’s theorem for nonnegative real coefficients), we get that indeed
$C_{h}(1)=\lim_{x\nearrow 1}C_{h}(x)=\lim_{x\nearrow 1}\big(1-\frac{1}{B_{h}(x)}\big)=1-\frac{1}{\lim_{x\nearrow 1}B_{h}(x)}=1-\frac{1}{B_{h}(1)}\,.$}
Here $\lim_{x\nearrow 1}$ denotes that $x$ approaches $1$ through the values $x\in[0,1)$.
Thus
$\lim_{n\to\infty}P_{d}(\overline{0},\,n)=\lim_{n\to\infty}h(A_{n})=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}h(C_{j})=C_{h}(1)=1-\frac{1}{B_{h}(1)}\,.}$
Since $B_{h}(1)=+\infty$ for $d\leq 2$ and $1<B_{h}(1)<+\infty$ for $d\geq 3$ (this requires some work), Pólya’s theorem for $\overline{v}=\overline{0}$ follows. $\Box$
See Theorem 3.5 for a generalization of this result. We have not seen the formula
$\lim_{n\to\infty}P_{d}(\overline{0},\,n)=\lim_{n\to\infty}h(A_{n})=1-\frac{1}{B_{h}(1)}\,$,} (2)
which unites the cases $d\leq 2$ and $d\geq 3$ of Pólya’s theorem for $\overline{v}=\overline{0}$, in the literature.
The use of generating functions in the proof of Pólya’s theorem is well known, but the theorem itself is mostly viewed as a result in probability or random walks; see, for example, [6] for a probabilistic proof. One of our goals is to convince the reader that Pólya’s theorem can also be naturally viewed as a result in enumerative combinatorics. The latter interpretation does not require tools of probability theory, such as Markov chains.
1.3 Pólya’s theorem for $\overline{v}\neq\overline{0}$ via the symbolic method
Let $\overline{v}\neq\overline{0}$ and $A_{\overline{v}}$ be the set of walks in $G_{d}$ starting at $\overline{0}$ and visiting $\overline{v}$. In the case $d\leq 2$ when $B_{h}(1)=+\infty$ we have, as a special case of item 1 in Theorem 4.13, that
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)=\lim_{n\to\infty}h(A_{\overline{v},\,n})=1\,.$ (3)
In the case $d\geq 3$ when $1<B_{h}(1)<+\infty$, we can obtain, by a more involved version of the argument in Section 1.2, the formula
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)=\lim_{n\to\infty}h(A_{\overline{v},\,n})={\textstyle\sqrt{1-\frac{B_{\overline{v},\,h}(1)}{B_{h}(1)}}}\,.$ (4)
Here $B_{\overline{v}}$ ($\subset B$) is the set of walks in $G_{d}$ that begin and end at $\overline{0}$ and avoid $\overline{v}$. It is easy to see that
$1\leq B_{\overline{v},\,h}(1)<B_{h}(1)\,$,
which implies the second case in (1). This result is generalized in Theorem 4.9. We have not seen the formula (4) in the literature.
1.4 General Pólya’s theorems
The idea of generalizing Pólya’s theorem is simple. If we extend the grid graph $G_{d}$ to the countable complete graph
$K=\big\langle\mathbb{Z}^{d},\,\binom{\mathbb{Z}^{d}}{2}\big\rangle\,$,}
extend the weight $h\colon E_{d}\to\{\frac{1}{2d}\}$ to the weight $h\colon\binom{\mathbb{Z}^{d}}{2}\to\{0,\frac{1}{2d}\}$ by setting $h=0$ on $\binom{\mathbb{Z}^{d}}{2}\setminus E_{d}$, if we then extend $h$ multiplicatively to walks in $K$ (see below), and if $A_{\overline{v}}$ is the set of walks in $K$ starting at $\overline{0}$ and later visiting the given vertex $\overline{v}$, then nothing changes in the above discussion of Pólya’s theorem in Sections 1.2 and 1.3. So we generalize it by allowing any complex weight $h\colon\binom{\mathbb{Z}^{d}}{2}\to\mathbb{C}$, assuming that a convergence condition holds so that the coefficients $h(A_{\overline{v},n})$ and coefficients in auxiliary generating functions, which are now sums of (possibly) infinite series, exist.
For simplicity we replace $\mathbb{Z}^{d}$ with the set $\mathbb{N}$ of natural numbers and work with the countable complete graph
$K_{\mathbb{N}}=\langle\mathbb{N},\,\mathbb{N}_{2}\rangle,\ \mathbb{N}_{2}=\binom{\mathbb{N}}{2}\,$.}
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight. $\mathbb{C}$ denotes the field of complex numbers, with the usual complete Archimedean norm (absolute value) $|z|=\sqrt{z\cdot\overline{z}}$. A walk $w$ in $K_{\mathbb{N}}$ is an $n+1$-tuple of vertices
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle,\,\text{$n\in\mathbb{N}_{0}$ and $u_{i}\in\mathbb{N}$}\,$,
such that $u_{i-1}\neq u_{i}$ for every $i=1,2,\dots,n$. Let $W$ be the set of all walks and $s\colon W\to\mathbb{N}_{0}$ be the size function given by $s(w)=n$, that is, $s(w)$ is the length of $w$. We standardly extend the weight $h$ to $W$: $h(w)=1$ if $s(w)=0$, and
$h(w)=\prod_{i=1}^{n}h(\{u_{i-1},\,u_{i}\})$}
if $s(w)=n\geq 1$. By the reversal of walks we mean the bijection (even involution) $\varphi\colon W\to W$,
$\varphi(\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle)=\langle u_{n},\,u_{n-1},\,\dots,\,u_{0}\rangle\,.$
The map $\varphi$ is length-preserving and weight-preserving: always $s(w)=s(\varphi(w))$ and $h(w)=h(\varphi(w))$.
For $n\in\mathbb{N}_{0}$ we denote by $D_{n}$ the set of walks of length (size) $n$ in $K_{\mathbb{N}}$ starting at $1$. Thus $D_{0}=\{\langle 1\rangle\}$ and for $n\geq 1$ the set $D_{n}$ is infinite and countable. Let $v\in\mathbb{N}$. We denote by $A_{v}$ the set of walks in $K_{\mathbb{N}}$ starting at $1$ and later visiting $v$. Then
$A_{v,n}=(A_{v})_{n}=\{\langle u_{0},\,\dots,\,u_{n}\rangle\in W\colon\;\text{$u_{0}=1$ and $\exists\,i>0$ such that $u_{i}=v$}\}\,.$
The said convergence condition is Definition 2.1: every series
$\sum_{w\in D_{n}}h(w),\ n\in\mathbb{N}_{0}\,$,}
absolutely converges. Absolute convergence is defined at the beginning of the next section. Then every series $\sum_{w\in A_{v,n}}h(w)$, $n\in\mathbb{N}_{0}$, absolutely converges as well; we denote its sum by $h(A_{v,n})$. We consider the main generating function
$A_{v,\,h}(x)=(A_{v})_{h}(x)=\sum_{n\geq 0}h(A_{v,\,n})x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
The classical theorem of Pólya is the formula (1). Many other graph families besides $G(d)$ have been investigated. What should the general Pólya’s theorems be about? It is not so clear, weights in them are general, $h\colon\mathbb{N}_{2}\to\mathbb{C}$, and the original probabilistic interpretation disappears in the distance. We investigate a variety of weights: general weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$, $\xi$-convex weights with $\xi\in\mathbb{C}$ and $|\xi|=1$, 0-convex weights (see Definition 2.10), weights $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ and weights $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$. Our answer is that general Pólya’s theorems are identities of the forms (2)–(4) for the walks $A_{v}$ in terms of quantities of four kinds, which we present below in the order of decreasing precision.
  1. 1.
    The limit $\lim_{n\to\infty}h(A_{v,n})$ of total weights of walks in $A_{v,n}$.
  2. 2.
    The sum $A_{v,h}(1)_{2}$ of absolutely convergent series
    $h(A_{v,0})+h(A_{v,1})+\dots$
    of those total weights.
  3. 3.
    The sum $A_{v,h}(1)_{1}$ of conditionally convergent series
    $h(A_{v,0})+h(A_{v,1})+\dots\$
    of those total weights.
  4. 4.
    The generalized sum
    $\lim_{x\nearrow 1}A_{v,h}(x)=\lim_{x\nearrow 1}{\textstyle\sum_{n\geq 0}h(A_{v,\,n})x^{n}}\,.$
In accordance with the ranges of weights, these quantities are usually twisted by powers of $\xi$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$). In Sections 3 and 4 we obtain along these lines a total of sixteen general Pólya’s theorems.
Some questions remain. In Theorem 3.2 we obtain the formula $A_{h}(1)_{1}=(1-\frac{1}{B(1)_{1}})D_{h}(1)_{1}$ which we deduce from the (simpler!) formula $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. Why do we prefer the former formula over the latter? First, the walks $B$ (starting and ending at $1$) have simpler definition than the walks $C$ (starting and ending at $1$ and avoiding $1$ between). Second, (1) follows from formulas of the former type, but it is not clear how to deduce it from the latter. See also the discussion after Proposition 4.3.
1.5 Plan of the article
In this section, we have hopefully explained the main ideas of our article and its context. To work with generating functions whose coefficients are sums of infinite series, we need a theory of complex series defined on arbitrary countable sets, such as the sets $D_{n}$ of walks in $K_{\mathbb{N}}$ of length $n$ starting at $1$. We could not find such a theory in the literature, at least not in the form we need, and therefore we develop it in Section 2. Here is one example. The product rule (formula) in the symbolic method, the identity that
$|A\times B|=|A|\cdot|B|$
for any two finite sets $A$ and $B$, is so simple and obvious that one encounters difficulties when asked to provide a detailed proof: first, axioms of set theory and natural numbers have to be specified, and so on. The actual product formula operates on the level of generating functions, ordinary or exponential ones, and is less immediate, but it has at its heart the displayed set-theoretic identity; see [3, pp. 22, 40], [8, Theorem on pp. 3–4], [11, Definition I.6 on p. 23] and [13, 2.2.14 Lemma on pp. 36–37]. In our extension of the symbolic method, we lift the product rule to the nontrivial Proposition 2.5 on products of absolutely convergent series. Proposition 2.4 concerns another important operation on series, the grouping. Other useful results in Section 2 are three Abel’s Theorems 2.162.18 (stated without proofs) and various conversions of formal identities to identities at $x=1$ (we saw one example in Section 1.2), such as Proposition 2.30.
In Section 3 we present generalizations of Pólya’s theorem in the simple case $v=1$ when the vertex to be visited by walks in $K_{\mathbb{N}}$ coincides with their starting vertex. In Proposition 3.1, we give detailed proofs of formal identities
$A_{h}(x)=C_{h}(x)D_{h}(x)\,\text{ and }\,C_{h}(x)=1-\frac{1}{B_{h}(x)}\ \ (\in\mathbb{C}[[x]])$}
whose instances for the graph $G_{d}$ we saw in Section 1.2. At first glance, they look like typical identities produced in the first phase of the symbolic method, but they are in fact completely new. The coefficients in the generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ arise as sums of infinite (absolutely convergent) series. In the symbolic method, such coefficients always arise as finite sums or finite linear combinations. The proof of Proposition 3.1 therefore transcends analogous proofs in the standard symbolic method. Our seven general Pólya’s theorems for $v=1$ are given in Theorems 3.23.8. As we already mentioned, these theorems show that formulas (2)–(4) hold for weights $h$ far more general than probabilistic ones and they explore the logical space of conditional and absolute convergence assumptions on the values at $x=1$ of generating functions. See Section 5 for the summary of these theorems.
Results in Section 4 in Theorems 4.54.13 are analogous to Section 3 but they are obtained in the more complicated case $v\neq 1$ of general Pólya’s theorems; see Section 5 for the summary of these theorems.. Proofs of the formal identities
$A_{v,\,h}(x)=C_{v,\,h}(x)D_{h}(x),\ B_{h}(x)=B_{v,\,h}(x)+C_{v,\,h}(x)^{2}B_{h}(x)\,\text{ and}$
$C_{v,\,h}(x)=B_{v,\,h}(x)E_{v,\,h}(x)$
in Proposition 4.2 are only sketched. For $v\neq 1$ one can introduce more auxiliary generating functions than for $v=1$, and in Proposition 4.3 we give another two formal identities for the generating function $F_{v,h}(x)$ of walks going from $1$ to $v$. In Theorem 4.4 we show that for any weight $h\colon\mathbb{N}_{2}\to\mathbb{R}$ and any vertex $v\neq 1$ we have, for every real $x\in(-1,1)$ and under appropriate assumptions, that if $B_{h}(x)=0$ then $B_{v,h}(x)=0$, and if $B_{h}(x)\neq 0$ then
$\frac{B_{v,\,h}(x)}{B_{h}(x)}\leq 1\,.$
This generalizes the trivial bound
$0\leq B_{v,\,h}(x)\leq B_{h}(x)$
for $x\in[0,1)$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$.
1.6 Perspectives and further research
We already mentioned that in [17] we develop a semiformal theory of formal power series. Many isolated instances of such uses of formal power series can already be found in the literature, and in [17] we want to systematize them. For example, the classical identity
$1-\frac{(\pi x)^{2}}{3!}+\frac{(\pi x)^{4}}{5!}-\dots=\prod_{n\geq 1}\Big(1-\frac{x^{2}}{n^{2}}\Big)\ \ (\in\mathbb{C}[[x]])$
resulting from the representation of the function $\sin(\pi x)/(\pi x)$ by an infinite product can be understood as an identity for formal power series employing semiformal operations. In [19] we generalize Pólya’s theorem to $K_{\mathbb{N}}$ with edge weights in the ring $\mathbb{C}[[x_{1},\dots,x_{k}]]$. There the norm is non-Archimedean, and we are curious what results we obtain. Weights with values in polynomial or formal power series rings are common in enumerative combinatorics. It would be interesting to obtain general Pólya’s theorems for several (maybe infinitely many?) vertices $v\in\mathbb{N}$ to be visited by walks in $K_{\mathbb{N}}$ starting at $1$.
@@ 2 Series and generating functions in $\mathbb{C}$
In this section, we develop a theory of general complex series and obtain propositions on the conversion of formal identities for generating functions to identities at $x=1$. We use these results in the two following sections.
2.1 General complex series
series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it also as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@ 2.1 General complex series
    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
@@
In fact, the independence of $s$ on $f$ may be omitted from the condition. We call the number $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$ or by $h(X)$. If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
@@
is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $U(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
@@
where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. So for any formal power series $U(x)\in H_{1}$ the sums $U(c)$ form the function
$U\colon\mathbb{C}_{<1}\to\mathbb{C},\ c\mapsto U(c)\,.$
Next definition is basic in our approach. Recall that $D_{n}$ is the set of walks in $K_{\mathbb{N}}$ with length $n$ starting at $1$.
Definition 2.1 (light weights)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight. We call $h$ light if for every $n\in\mathbb{N}_{0}$, the series
$\sum_{w\in D_{n}}h(w)$}
@@
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare the sizes of sums $h(Y)$ and $h(X)$; we may have $h(X)=0$ and $h(Y)\neq 0$. However, if $h\colon X\to\mathbb{R}_{\geq 0}$, $\sum_{x\in X}h(x)$ absolutely converges, and if $Y\subset X$, then obviously
$0\leq h(Y)\leq h(X)\,.$
We call it the subseries bound.
2.2 Linear combination, grouping and product of series
Suppose that $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and that $\alpha,\beta\in\mathbb{C}$. Then the series